OpenStudy (anonymous):

Solve differential: 8y' = x+y, change of variable u = x + y Medal for showing me the solutions! I got y = e^(1/8x)e^C-8-x These confuse me, please help. If I just see how I can do it

OpenStudy (anonymous):

To me, I would solve it as 8y'-y =x for homogenous part, the characteristic equation is 8r-1=0 --> r = 1/8 so the general solution is y = Ce^(1/8x) for partial, y = Ax +B y' = A so 8y'-y= 8A -Ax -B = x --> A = -1 and B = -8 so partial solution is -x-8 combine them I get \(y = Ce^{x/8}-x-8\)

OpenStudy (anonymous):

we're covering "separable equations" is that what you mean by partial? so C is in place for e^C right? if that's the case we have the same answer, why is it wrong without the constant C before e?

OpenStudy (anonymous):

nope not separable. It is the form of y'+P(x) y = Q(x) where P(x) = 1 and Q(x) = x

OpenStudy (anonymous):

I just don't understand why C must be a dummy variable for e^C? but thank you for helping :)

OpenStudy (anonymous):

that's the formula. I don't know why C but not e^C. however, e^C = C, they are the same because they are some constant.

OpenStudy (anonymous):

ok, I understand that. THank you again! :)